OpenStudy (abbster2015backup):

Suppose p varies directly with d, and p = 7 when d = 3. What is the value of d when p = 28? A. 140/3 B. 4/3 C. 12 D. 16

OpenStudy (abbster2015backup):

@MathMusician

OpenStudy (mathmusician):

one sec

OpenStudy (mathmusician):

If p varies directly with d, then p = k * d 7 = k * 3, so k = 7/3 Now, use that value to find the value of d. You are also given that p = 28 in this case. 28 = (7/3)d ... now just solve this equation for d. 7d = 28 * 3 7d = 84 d = 12

OpenStudy (abbster2015backup):

Oh so it's C?

OpenStudy (abbster2015backup):

@PillowFace750 @ganeshie8

OpenStudy (abbster2015backup):

@andriod09

OpenStudy (andriod09):

You already have the answer, hon. Why're you tagging others? The question asked wonders what the value of "d" is, when "d" relates to "p", and as @Mathmusician has already explained, "d=12" would be the answer. :)

OpenStudy (abbster2015backup):

Oh, I thought that was something else xDD So it's C, thanks!

OpenStudy (andriod09):

No problem! Don't forget to hand out a medal to who you felt helped you the most and also to close the question if you're done. :)